# Prompt Stellar and Binary Black Hole Mergers from Chemically Homogeneous Evolution in Triples

### v1

@@ Acknowledgements.
The authors thank Sunmyon Chon, Ilya Mandel, Silvia Toonen, and Ruggero Valli for helpful discussions. EG and AB acknowledge support from the Australian Research Council (ARC) Centre of Excellence for Gravitational Wave Discovery (OzGrav), through project number CE230100016. EG further acknowledges support from the ARC Discovery Program DP240103174 (PI: Heger). This research made use of MESA: Modules for Experiments in Stellar Astrophysics, r11701 (http://mesa.sourceforge.net, Paxton et al., 2011, 2013, 2015, 2018, 2019). The MESA inlists used used to model binary evolution are available via zenodo, with DOI 10.5281/zenodo.3667546 (du Buisson et al., 2019). The scripts used for the dynamical analyses in this study are available on GitHub at avigna/che-triples. Kippenhahn diagram plotter for MESA available in GitHub at orlox/mkipp.
@@
In this paper, we use detailed stellar evolution to explore the chemically homogeneous evolution (CHE) of a binary (§2.1.1). To do so, we follow Sharpe et al. (2024), which focuses on HD 5980, a multiple-star system located within the Small Magellanic Cloud. To model HD 5980, Sharpe et al. (2024) adopts a default metallicity of $Z=Z_{\odot}/4=0.00425$, using $Z_{\odot}=0.017$ (Grevesse et al. 1996). However, they acknowledge that this value is twice as high as the commonly accepted metallicity for the Small Magellanic Cloud. Given that CHE is highly sensitive to metallicity (e.g., Marchant et al. 2017; Riley et al. 2021), we explore an alternative model with higher metallicity while maintaining all other assumptions.
[…]
@@
Figure 9: Same as Fig. 1 but for $Z=0.0042$. The blue dashed vertical line here is shown at at 5.31 Myr.

### v2

@@ Acknowledgements.
The authors thank Sunmyon Chon, Selma de Mink, Ilya Mandel, Luca Sciarini, Silvia Toonen, and Ruggero Valli for helpful discussions. We thank the referee Ying Qin for useful suggestions. EG and AB acknowledge support from the Australian Research Council (ARC) Centre of Excellence for Gravitational Wave Discovery (OzGrav), through project number CE230100016. EG further acknowledges support from the ARC Discovery Program DP240103174 (PI: Heger). BL acknowledges support from National Natural Science Foundation of China (Grant No. 12433008). This research made use of MESA: Modules for Experiments in Stellar Astrophysics, r11701 (http://mesa.sourceforge.net, Paxton et al., 2011, 2013, 2015, 2018, 2019). The MESA inlists used used to model binary evolution are available via zenodo, with DOI 10.5281/zenodo.3667546 (du Buisson et al., 2019). The scripts used for the dynamical analyses in this study are available on GitHub at avigna/che-triples. Kippenhahn diagram plotter for MESA available in GitHub at orlox/mkipp.
@@
In this paper, we use detailed stellar evolution to explore the CHE of a binary (Sect. 2.1.1). To do so, we followed the study of Sharpe et al. (2024), focused on HD 5980, a multiple-star system located within the SMC. To model HD 5980, Sharpe et al. (2024) adopts a default metallicity of $Z=Z_{\odot}/4=0.00425$, using $Z_{\odot}=0.017$ (Grevesse et al. 1996). However, they acknowledge that this value is twice as high as the commonly accepted metallicity for the SMC. Given that CHE is highly sensitive to metallicity (e.g. Marchant et al. 2017; Riley et al. 2021), we explore an alternative model with higher metallicity while maintaining all other assumptions.
In our main analysis, we focussed on a default CHE binary at $Z=0.00042$ (Sect. 2.1), which is 10% of the default value used in Sharpe et al. (2024). Here, we present the results of our CHE binary model at $Z=0.0042$. Figure 9 illustrates the stellar structure. At the ZAMS, the total mass and the mass of the convective core are similar to those at the default metallicity. However, both the radius of the star ($\approx 10R_{\odot}$) and the convective core ($\approx 4R_{\odot}$) are slightly larger,which is also reflected in a lower value for the apsidal motion constant. Throughout its evolution, this stellar model loses significantly more mass compared to the default model, finishing its evolution with $\approx 18.6M_{\odot}$. This is a direct consequence of line-driven winds, which are more efficient at higher metallicities. Mass loss through winds leads to a widening of the orbital period to $\approx 4.5$ d. If such a CHE binary were isolated and became a BBH, with negligible mass loss and change in the orbital period, it would take 25.5 Gyr to merge via gravitational-wave emission.
@@
Figure 9: Same as Fig. 1 but for $Z=0.0042$. The vertical dashed blue line represents 5.31 Myr.